OpenStudy (cenaida):

Use the Pythagorean Theorem to find the length of the hypotenuse in a right triangle whose legs are 3 and 4. A. 3 B. 5 C. 4 D. 2

OpenStudy (anonymous):

Okay, so in order to get your answer, all you have to do is plug in the information you are given into the Pythagorean theorem, a^2+b^2=c^2.

OpenStudy (anonymous):

a=3 b=4 3^2=9 4^2=16 16+9=25 sqrt of 25 is 5 so 5^2. Your Answer is B. 5

OpenStudy (anonymous):

Since its a Pythagorean triple though, you don't even have to do that. you know its 5.

OpenStudy (anonymous):

Just for future reference.

OpenStudy (cenaida):

Oh my god thank you guys! so much!!!!

OpenStudy (cenaida):

so is it the same for this question? What is the length of the leg in a right triangle whose hypotenuse is 13 and other leg is 12. A. 3 B. 5 C. 4 D. 2

OpenStudy (anonymous):

so a^2 +12^2 = 13^2 a^2 + 144 = 169 a^2 (-144) = 169 -144 a^2 = 25 a = √25 a = 5 so you're answer is 5.